OpenStudy (anonymous):

Please help me figure out how to create a quadratic equation in standard form that can be factored.

OpenStudy (cwrw238):

well one way ss to think of 2 numbers a and b where ab = say 6, now find some factors of 6 for example 3 and 2 so let a = 3 and b = 2 (x +a)(x + b) = 0 (x + 3)(x + 2) = 0 expand to get x^2 + 5x + 6 = 0

OpenStudy (cwrw238):

what we are doing is finding integers to fit the expression x^2 + ax + bx + ab

OpenStudy (cwrw238):

we can of course have negatives fo example x^2 - 2x + 7x - 14

OpenStudy (cwrw238):

giving x^2 + 5x - 14 = 0

OpenStudy (anonymous):

honestly I'm very confused I think we should start at the top and maybe you could explain this a little easier to me.

OpenStudy (cwrw238):

Ok the factored form of a trinomial ( or quadratic) as like (x + a)(x + b) = 0 now if we expand this we get x^2 + ax + bx + ab ok?

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

so we want to fit integers to a and b lets pick ab ( a times b) first ab = 8 now we can choose 2 integers that when multiplied give 8 we can pick 2 and 4 , or if we like 8 and 1 lets pick 2 and 4 now plug these in to the expression x^2 + ax + bs + ab = x^2 + 2x + 4x + 8 = x^2 + 6x + 8 so our quadratic equation is x^2 + 6x + 8 = 0

OpenStudy (cwrw238):

just keep in mind the identity x^2 + ax + bx + ab and its easy

OpenStudy (anonymous):

so is that equation in standard form?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

oh ok so how would we put it into factored form

OpenStudy (cwrw238):

well x^2 + 6x + 8 = 0 is in standard form its more accurate to write the a b form as x^2 + (a+ b)x + ab = 0 standard form

OpenStudy (cwrw238):

to facto x^2 + 6x + 8 = 0 by observation we see that 2*4 = 8 and 2 + 4 = 6 so we write x^2 + 2x + 4x + 8 = 0 now factor the first 2 terms and the last 2 terms ( this is called grouping) x(x + 2) + 4(x + 2) = 0 note that the parentheses are equal so we can write (x+ 2)(x + 4) = 0